OpenStudy (anonymous):

could anybody help me please simplify the rational expression 3x-6/X-2

OpenStudy (anonymous):

\[\frac{3x-6}{x-2}=\frac{3(x-2)}{x-2}=3\]

OpenStudy (anonymous):

as long as \(x\neq 2\)

OpenStudy (anonymous):

so the answer is 3? could you help me with another please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

SIMPLFY RADICAL 1.square root 7( squareroot 14 + squareroot 3) 2.square root 75+square root 3

OpenStudy (anonymous):

\[\large \sqrt{7}(\sqrt{14}+\sqrt{3})\] \[=\large \(\sqrt{7\times 14}+\sq

OpenStudy (anonymous):

\[\large =\sqrt{7\times 14}+\sqrt{7\times 3}\] \[\large=\sqrt{98}+\sqrt{21}\] \[=\sqrt{49\times 2}+\sqrt{21}\] \[=\sqrt{49}\sqrt{2}+\sqrt{21}\] \[=7\sqrt{2}+\sqrt{21}\]

OpenStudy (anonymous):

\[\sqrt{75}+\sqrt{3}=\sqrt{25\times 3}=\sqrt{3}=5\sqrt{3}+\sqrt{3}=6\sqrt{3}\]

OpenStudy (anonymous):

thank you can you help me with another